# Structured Sparse Transition Matrices to Enable State Tracking in State-Space Models

### v2

@@ B.2 Efficiently Computing the Parallel Scan Gradients
$\displaystyle\frac{dL}{\partial x_{k}}=\frac{\partial L}{\partial x_{k}}+\frac{dL}{dx_{k+1}}A_{k+1}$
@@ Appendix C Proofs
Property 3 (Algebraic Structure of One-hot Column Matrices).
@@
Property 4 (Computational Efficiency of Matrix Multiplication in $\mathbb{H}^{N\times N}$).
@@
$\mathbb{V}:=\mathrm{span}\{v_{1},\ldots,v_{N}\}=\mathrm{span}\{v_{1},\ldots,v_{N-2}\}\text{ and }v_{N-1}=\sum_{i=1}^{N-2}\alpha_{i}v_{i}\text{ as well as }v_{N}=\sum_{i=1}^{N-2}\beta_{i}v_{i}.$
@@
According to Jaffe [2007], $k_{t}$ is linearly independent over $\mathbb{Q}$, i.e., for $u_{i}\in\mathbb{Q}$ it holds that $u_{1}k_{1}+\cdots+u_{n}k_{n}=0\iff u_{1}=\cdots=u_{n}=0.$ Therefore, no two differing input sequences $u_{1},\ldots,u_{n}$ and $v_{1},\ldots,v_{m}$ will result in the same state representation $\textstyle x_{t}=x_{0}+\sum_{i}u_{i}k_{i}$. Now, a simple identification of every FSA state with all the state representations $x_{t}$ of associated input sequences that lead to said FSA state shows universal expressivity under arbitrary readout. ∎
@@ D.1 Hyperparameter Selection
On FSA emulation, Table 2, we did not perform a hyperparameter grid search. We re-used the fixed hyperparameters which were used to evaluate all of the baseline methods, as per Walker et al. [2025]. In contrast to the baseline methods which were evaluated using various state sizes including 128, we only evaluated our model using state size $128$.
On experiments with non-solvable groups in Table 3, we ran a learning rate grid search in $\{1e{-4},5e{-4},1e{-3},5e{-3},1e{-2}\}$, and report the best validation accuracy over any of the 3 random seeds and hyperparameter configurations. The state size was 128.
@@
We performed the following grid search, following Rusch and Rus [2025]: Learning rate in $\{1e{-5},1e{-4},1e{-3}\}$, number of layers in $\{2,4,6\}$, state dimension in $\{16,64,128\}$ and embedding dimension in $\{16,64,128\}$.
@@
We performed a grid search defined by all combinations of the following values: Learning rate in $\{1e{-5},5e{-5},1e{-4},5e{-4},1e{-3}\}$, weight decay in $\{0.001,0.01,0.1,0.05\}$, and dropout in $\{0.0,0.1,0.3\}$. On each task, we report the average accuracy over three random seeds. The embedding dimension and state size were both fixed to 128, with the exception of the Retrieval dataset in which we reduced it to 64. We used Adam with the default parameters (0.9, 0.999). The hyperparameters of the models we compare with were obtained via an extensive Bayesian hyperparameter search, as outlined in Soydan et al. [2024].
@@ Appendix E Additional Results
E.1 Runtime Experiments
@@
The runtime measurements were obtained using optimized and compiled JAX code [^5] https://docs.jax.dev/en/latest/quickstart.html. We did not implement custom system-aware methods, which can certainly make a difference in practice [Gu and Dao 2023, Dao 2024]. We assume that our runtime measurement setup is reflective of most researchers’ typical use cases, and should provide an equal testbed for all the methods. All of the runtime measurements were obtained on an NVIDIA A100 80GB GPU. The code is implemented in JAX, version 0.4.24. The parallel scan relies on the jax.lax.associative_scan primitive. The computation is parallelized across the batch dimension and sequence length using jax.vmap, and is finally just-in-time compiled. At the time of writing, the parallel scan primitive is a work-in-progress in PyTorch. In our experience, the JAX version was significantly more efficient.
Runtime vs. Sequence Length
In this section, we will compare our PD model with two representative methods, a time-variant SSM with unstructured but column-normalized transition matrices [^6] Column normalization is a crucial step to ensure system stability when using unstructured transition matrices, and is typically implemented by dividing the matrix elements by the $L^{p}$ quasi-norm of the corresponding column [Fan et al. 2024, Terzić et al. 2025]. As a consequence of the Gershgorin disc theorem and basic inequalities, a matrix with $L^{p}$-normalized columns has eigenvalues smaller than 1 when $p<1$., and a time-variant SSM with diagonal transition matrices generated using our MLP generation. We report the runtime as a function of sequence length for the three models under consideration, using three different state sizes, 32, 128 and 1024. For both state sizes, we increase the batch size to the largest power of 2 which does not induce an out-of-memory error with the longest sequence. Results are shown in Tables S1S2, and S3. The benefit of our method increases with state size. In the following section, we identify the main bottleneck of our method not as the parallel scan, but rather as the $P$ matrix generation.
Table S1: Runtimes [s] with state size $N$= 32, batch size $B$= 1024
Sequence Length 128 640 1152 1664 2176 2688 3200 3712 4224 Unstructured SSM 0.00621 0.03074 0.05520 0.07962 0.10456 0.12933 0.15392 0.17828 0.20330 PD-SSM 0.00377 0.01749 0.03106 0.04724 0.06148 0.07457 0.08989 0.10331 0.11638 Diagonal SSM 0.00174 0.00554 0.00934 0.01366 0.01853 0.02203 0.02638 0.03037 0.03528
Table S2: Runtimes [s] with state size $N$= 128, batch size $B$= 64
Sequence Length 128 640 1152 1664 2176 2688 3200 3712 4224 Unstructured SSM 0.00429 0.02089 0.03755 0.05551 0.07121 0.08776 0.10446 0.12104 0.13806 PD-SSM 0.00189 0.00827 0.01427 0.02172 0.02683 0.03295 0.03908 0.04520 0.05159 Diagonal SSM 0.00044 0.00165 0.00293 0.00285 0.00389 0.00440 0.00517 0.00587 0.00652
Table S3: Runtimes [s] with state size $N$= 1024, batch size $B$= 1
Sequence Length 128 640 1152 1664 2176 2688 3200 3712 4224 Unstructured SSM 0.00625 0.02997 0.05402 0.07811 0.10255 0.12690 0.15140 0.17588 0.19948 PD-SSM 0.00136 0.00492 0.00846 0.01194 0.01810 0.02232 0.02644 0.03065 0.03469 Diagonal SSM 0.00024 0.00033 0.00053 0.00061 0.00074 0.00097 0.00090 0.00156 0.00173
Effect of $P$ Generation on Runtime
@@
In this measurement, we compare the runtime of the PD-SSM with a version of PD-SSM that receives pre-computed random sparse $P$ matrices as input. The results are reported in Table S4.
Table S4: Runtimes [s] with state size $N$= 1024, batch size $B$= 1
Sequence Length 128 640 1152 1664 2176 2688 3200 3712 4224 Unstructured SSM 0.00625 0.02997 0.05402 0.07811 0.10255 0.12690 0.15140 0.17588 0.19948 PD-SSM (full) 0.00136 0.00492 0.00846 0.01194 0.01810 0.02232 0.02644 0.03065 0.03469 PD-SSM (Random $P(u_{t})$) 0.00045 0.00084 0.00097 0.00146 0.00185 0.00212 0.00241 0.00275 0.00258 Diagonal SSM 0.00027 0.00041 0.00054 0.00060 0.00090 0.00086 0.00113 0.00159 0.00170
E.2 Varying Mamba Depth on $A_{5}$
@@
Mamba [Gu and Dao 2023] uses non-negative real-valued transition matrices, which are in theory highly restrictive even for emulating solvable automata [Grazzi et al. 2024] [^7] Such models must converge to a steady state upon application of a constant input. This goes contrary to FSA emulation. For example, Parity never converges upon repeated application of a 1.. However, we can observe that the model does in fact learn to emulate the non-solvable automaton on in-domain lengths, as shown in Figure S10. It does, however, not exhibit any substantial length generalization.
@@
Figure S10: Mamba with varying number of layers trained to emulate the non-solvable $A_{5}$ automaton.
E.3 Additional State Tracking Results
This set of experiments is based on the setup from Delétang et al. [2023] and measures the ability of a model to track the states of FSAs. The one-hot input symbols are mapped to an embedding through a trainable dictionary.We use two FSAs from Delétang et al. [2023], namely Parity and Arithmetic. Furthermore, we define three FSAs with group structure; $\mathbb{Z}_{2}\times\mathbb{Z}_{30}$ is a commutative, solvable automaton, $D_{30}$ is a non-commutative solvable automaton, and $A_{5}$ is a non-solvable automaton. The models are trained to predict the final state of the automaton given a sequence of inputs, with no
supervision on intermediate states. During training, random sequences of length up to 40 are sampled. During testing, the models are evaluated on sequences of length up to 500. Both Mamba [Gu and Dao 2023] and our complex diagonal SSM defined by omitting the $P$ matrices struggle on non-solvable automata, confirming their theoretical expressivity summarized in Table 1. In contrast, SD-SSMs [Terzić et al. 2025] and PD-SSMs excel at state tracking, with the latter being much more eficient. We provide fur-
Task Mamba $\mathbb{C}$ Diag. SD-SSM PD-SSM
$C_{2}\times C_{30}$ 24.1 100 100 100
$D_{30}$ 18.9 61.0 100 100
$A_{5}$ 10.6 8.3 100 100
Table S5: Length generalization on FSA emulation tasks using single-layer SSMs. We report the average accuracy (%) over test sequences of length $1$ to $500$ for the best model from three random seeds of training.
ther ablations on the Arithmetic task in Appendix E.3, which demonstrate that (1) our deterministic approximation of $\partial P/\partial M$ outperforms a stochastic Gumbel-softmax straight-through estimator and that (2) the model achieves significantly better length generalization when using a linear readout function $\psi(\cdot)$
E.4 PD-SSM Ablations on Arithmetic
@@
We ablate several design choices on the Arithmetic task from Delétang et al. [2023].
@@
The dependence on $u_{t}$ as well as the indices are omitted for notational simplicity. Typically, straight-through gradient approximators, an example of which is our approximation of the hardmax via a softmax, are used in conjunction with categorical sampling schemes [Paulus et al. 2021, Jang et al. 2017, Bengio et al. 2013]. In particular, a widely used stochastic scheme takes the following form, called Gumbel max:
@@
with $||$ denoting vector concatenation. Typically, however, neural networks interleave sequence processing layers with nonlinearly activated feed-forward networks [Vaswani et al. 2017], as opposed to the affine transformation that we applied. We test the effect of a nonlinear readout, defined as
@@ Appendix F Related Work
LTI SSMs have been proposed as a more scalable alternative to Transformers that is also more performant on long sequences Gu et al. [2020, 2021, 2022b, 2022a], Gupta et al. [2022], Smith et al. [2023], Orvieto et al. [2023]. A subsequent analysis showed promising language modeling performance with hybrid architectures based on gated SSMs and Attention Ma et al. [2023], Fu et al. [2023]. Time-variant hybrid systems were then shown to perform significantly better on practical tasks with the Mamba SSM which utilizes diagonal transition matrices Gu and Dao [2023]. Consequently, a line of work has demonstrated that hybrid models utilizing diagonal time-variant SSMs and attention exhibit better generalization to longer sequences, better general language modeling and reasoning performance [Ren et al. 2025, De et al. 2024, Lenz et al. 2025, Wu et al. 2025, Waleffe et al. 2024].
@@
Concurrently, formal analyses of time-varying SSMs have shown limitations of the diagonal transition matrix structure Merrill et al. [2024], Cirone et al. [2024], spurring a line of work investigating the utility of block-diagonal and unstructured matrices Fan et al. [2024], Terzić et al. [2025], as well as more expressive structured transition matrices, most notably diagonal plus low-rank Yang et al. [2024], Grazzi et al. [2024], Peng et al. [2025], Siems et al. [2025], with the original inspiration stemming from at least as far back as fast-weight programmers Schlag et al. [2021a, b], Schmidhuber [1992] whose computational structure contains products of generalized Householder matrices Yang et al. [2024], which enables the adoption of early algorithmic optimizations of such procedures Bischof and Van Loan [1987].
@@
As a lower bound on diagonal SSMs, Sarrof et al. [2024] show that a stack of complex diagonal SSM layers can emulate any solvable automaton, with the stack depth proportional to the Krohn-Rhodes (KR) complexity of the transformation semigroup Krohn and Rhodes [1965], Margolis et al. [2024]. As an upper bound on diagonal SSMs, Merrill et al. [2024] derives a bound on fixed-depth diagonal selective SSMs with logarithmic precision representation, placing them in the $\text{L-uniform }TC^{0}$ circuit complexity class. This complexity class encompasses solvable automata. DPLR matrices can emulate any automaton with depth proportional to KR-complexity and width linear in depth and arbitrary readout Siems et al. [2025].
@@
Experimental analysis of sequential neural networks trained on formal language tasks goes back to at least [Smith and Zipser 1989, Hochreiter and Schmidhuber 1997]. The exact experimental setup we build on is a more recent one, with the set of tasks taken from Delétang et al. [2023], and the experimental setup borrowed from Walker et al. [2025]. All of the non-solvable group word problems were generated using modified sample generation code from Liu et al. [2023].

### v3

@@ B.2 Efficiently Computing the Parallel Scan Gradients
$\displaystyle\frac{dL}{dx_{k}}=\frac{\partial L}{\partial x_{k}}+\frac{dL}{dx_{k+1}}A_{k+1}$
@@ Appendix C Proofs
Property 1 (Algebraic Structure of One-hot Column Matrices).
@@
$\mathbb{V}:=\mathrm{span}\{v_{1},\ldots,v_{N}\}=\mathrm{span}\{v_{1},\ldots,v_{N-1}\}\text{ and }v_{N-1}=\sum_{i=1}^{N-2}\alpha_{i}v_{i}\text{ as well as }v_{N}=\sum_{i=1}^{N-2}\beta_{i}v_{i}.$
@@
According to jaffe2007linearly, $k_{t}$ is linearly independent over $\mathbb{Q}$, i.e., for $u_{i}\in\mathbb{Q}$ it holds that $u_{1}k_{1}+\cdots+u_{n}k_{n}=0\iff u_{1}=\cdots=u_{n}=0.$ Therefore, no two differing input sequences $u_{1},\ldots,u_{n}$ and $v_{1},\ldots,v_{m}$ will result in the same state representation $\textstyle x_{t}=x_{0}+\sum_{i}u_{i}k_{i}$. Now, a simple identification of every FSA state with all the state representations $x_{t}$ of associated input sequences that lead to said FSA state shows universal expressivity under arbitrary readout. ∎
@@ D.1 Hyperparameter Selection
On FSA emulation, Table 2, we did not perform a hyperparameter grid search. We re-used the fixed hyperparameters which were used to evaluate all of the baseline methods, as per walker_2025_structured. In contrast to the baseline methods which were evaluated using various state sizes including 128, we only evaluated our model using only state size $128$. We used Adam with the default parameters (0.9, 0.999).
On experiments with non-solvable groups in Table 3, we ran a learning rate grid search in $\{1e{-4},5e{-4},1e{-3},5e{-3},1e{-2}\}$, and report the best validation accuracy over any of the 3 random seeds and hyperparameter configurations. The state size was 128. We used Adam with the default parameters (0.9, 0.999).
@@
We performed the following grid search, following rusch2025oscillatory: Learning rate in $\{1e{-5},1e{-4},1e{-3}\}$, number of layers in $\{2,4,6\}$, state dimension in $\{16,64,128\}$ and embedding dimension in $\{16,64,128\}$. We used Adam with the default parameters (0.9, 0.999).
@@
We performed a grid search defined by all combinations of the following values: Learning rate in $\{1e{-5},5e{-5},1e{-4},5e{-4},1e{-3}\}$, weight decay in $\{0.001,0.01,0.1,0.05\}$, and dropout in $\{0.0,0.1,0.3\}$. On each task, we report the average accuracy over three random seeds. The embedding dimension and state size were both fixed to 128, with the exception of the Retrieval dataset in which we reduced it to 64. We used Adam with the default parameters (0.9, 0.999). The hyperparameters of the models we compare with were obtained via an extensive Bayesian hyperparameter search, as outlined in soydan_2024_simplified.
@@
D.2 Training Times
We report the training time for representative benchmarks using NVIDIA A100 40GB GPUS.
Natural Language State Tracking
This benchmark uses an inefficient PyTorch implementation of the PD-SSM, resulting in extended training times. The experiment can be executed significantly more efficiently using the provided JAX implementation.
On the Parity automaton:
  • PD : 7.2 hours
  • Complex diagonal: 70.3 hours
  • Real diagonal: 68.6 hours
On the $A_{5}$ automaton:
  • PD : 82.8 hours
  • Complex diagonal: 68.4 hours
  • Real diagonal: 63.6 hours
Long-Range Arena
  • CIFAR: 2.5 hours
  • ListOps: 9.3 hours
  • IMDB: 6.6 hours
  • AAN: 34 hours
UEA Time-Series
All results can be obtained within several hours on the specified GPU.
@@ Appendix E Additional Results
E.1 Runtime Measurements
Measurement Details
@@ E.1 Runtime Measurements
The runtime measurements were obtained using optimized and compiled JAX code [^5] https://docs.jax.dev/en/latest/quickstart.html. We did not implement custom system-aware methods, which can certainly make a difference in practice [gu_mamba_2023, dao2023flashattention2]. We assume that our runtime measurement setup is reflective of most researchers’ typical use cases, and should provide an equal testbed for all the methods. All of the runtime measurements were obtained on an NVIDIA A100 80GB GPU. The code is implemented in JAX, version 0.4.24. The parallel scan relies on the jax.lax.associative_scan primitive. The computation is parallelized across the batch dimension and sequence length using jax.vmap, and is finally just-in-time compiled. At the time of writing, the parallel scan primitive is a work-in-progress in PyTorch. In our experience, the JAX version was significantly more efficient.
State Representation Size Matched Comparison
We extend the comparison of our PD model with two representative methods, a time-variant SSM with unstructured but column-normalized transition matrices (SD-SSM) terzic2025sdssm  [^6] Column normalization is a crucial step to ensure system stability when using unstructured transition matrices, and has been implemented by dividing the matrix elements by the $L^{p}$ (quasi-)norm of the corresponding column [fan_advancing_2024, terzic2025sdssm]. As a consequence of the Gershgorin disc theorem and basic inequalities, a matrix with $L^{p}$-normalized columns is guaranteed to have spectral radius less than 1 whenever $p<1$., and a time-variant SSM with diagonal transition matrices generated using our MLP generation $A(u_{t})=D(u_{t})$. Below, we extend the results from Figure 4 by considering sequence lengths $256$ and $512$. Concretely, the dimension of SD-SSM is again set to twice that of the other models to account for the discrepancy between the real and complex-valued state representations. The results are shown in Figure S10.
(a) Sequence Length 256
(b) Sequence Length 512
Figure S10: Runtimes with state dimension of SD-SSM set to twice that of other models.
From Figure 4, at sequence length 64 and embedding dimension 5632, the SD-SSM exhibits an around $71\times$ slowdown compared to PD-SSM. The PD-SSM in turn is around $7\times$ slower than the diagonal model.
At hidden dimension $D=2048$, the relative speed-up of PD-SSM over SD-SSM is $18.8\times$ for $L=64$, $28.3\times$ for $L=256$, and $30.3\times$ for $L=512$, an increasing trend in $L$.
State Size Matched Comparison
Previously, the unstructured SSM had double the state dimensionality of the other two models to account for the discrepancy between the real and complex-valued state representations. We now set the state dimensionalities of all models to be equal and again evaluate the scaling as a function of embedding size under sequence lengths $64$ and $256$. The results are shown in Figure S11.
(a) Sequence Length 64
(b) Sequence Length 256
Figure S11: Runtime measurements with equal state size.
At sequence length 64 and embedding dimension 5632, the SD-SSM exhibits a $16.6\times$ slowdown compared to PD-SSM. The PD-SSM is in turn $7.6\times$ slower than the diagonal model, a similar ratio as in the state matched setup above but now with double the state size.
At hidden dimension $D=2048$, the relative speed-up of PD-SSM over SD-SSM is $6.8\times$ for $L=64$, $8.8\times$ for $L=256$, and $9.7\times$ for $L=512$ (not shown). Again, an increasing trend with $L$.
Parameter Matched Comparison
Given state size $N$, embedding dimension $D$ and $K$ transition matrices in the dictionary where applicable, the models have the following total parameter cost: $N(2D+KN)+KD$ for SD-SSM, $N(6D+2N+KN+4)+KD$ for PD-SSM, and $N(6D+2N+4)$ for the diagonal SSM. Let us denote the state size of SD-SSM as $N_{r}$ and that of PD-SSM as $N_{c}$. Suppose further that $D=N_{r}$ and $K=6$, as used in our measurements. To equalize the parameter cost of PD-SSM with SD-SSM, $N_{c}$ must be set as:
$N_{c}=\left\lfloor\frac{-(6N_{r}+4)+\sqrt{(6N_{r}+4)^{2}+192N_{r}^{2}}}{12}\right\rfloor$
With this, given for example $N_{r}=D=2048$, the corresponding $N_{c}$ is $1552$. The corresponding relationship for the state dimensionality of the diagonal model $N_{d}$ given $D=N_{r}$ and $K=6$ is:
$N_{d}=\left\lfloor\frac{-(6N_{r}+4)+\sqrt{(6N_{r}+4)^{2}+64N_{r}^{2}+48N_{r}}}{4}\right\rfloor$
which for $N_{r}=2048$ exactly sets $N_{d}=2048$. This adapted scaling results in the Figure S12.
(a) Sequence Length 64
(b) Sequence Length 256
Figure S12: Runtime measurements in a parameter-matched setting.
At sequence length 64 and embedding dimension 5632, the SD-SSM exhibits a $21.7\times$ slowdown compared to PD-SSM. The PD-SSM is now around $6.0\times$ slower than the diagonal model.
At hidden dimension $D=2048$, the relative speed-up of PD-SSM over SD-SSM is $10.6\times$ for $L=64$, $13.9\times$ for $L=256$, and $15.5\times$ for $L=512$ (not shown). Once again, an increasing trend with $L$.
Effect of $P$ Generation Under Equal State Size
@@
In this measurement, we compare the runtime of the PD-SSM and the diagonal SSM with a version of PD-SSM that receives pre-computed random sparse $P$ matrices as input. In this scenario, we set the state size of all models to be equal. The results are reported in Figure S13.
(a) Runtime comparison with length $L=64$
(b) Slowdown w.r.t. the diagonal baseline, $L=64$.
Figure S13: Measuring the overhead of generating $P$ matrices.
As predicted in Table 1, the $PD$ parallel scan incurs a constant, even diminishing overhead compared to the diagonal scan. The overhead stems primarily from the generation of the $P(u_{t})$ matrices.
E.2 Varying Mamba Depth on $(A_{5},2)$
@@
Mamba [gu_mamba_2023] uses non-negative real-valued transition matrices, which are in theory highly restrictive even for emulating solvable automata [grazzi2024unlocking] [^7] Such models must converge to a steady state upon application of a constant input. This goes contrary to FSA emulation. For example, Parity never converges upon repeated application of a 1.. However, we can observe that several layers of the model can in fact learn to emulate the non-solvable automaton on in-domain lengths, as shown in Figure S14. The model however does not exhibit any substantial length generalization.
@@
Figure S14: Mamba with varying number of layers trained to emulate the non-solvable $(A_{5},2)$ automaton.
E.3 PD-SSM Ablations on Arithmetic
@@
We ablate several design choices on the Arithmetic task from deletang_neural_2023.
@@
The dependence on $u_{t}$ as well as the indices are omitted for notational simplicity. Typically, straight-through gradient approximators, an example of which is our approximation of the hardmax via a softmax, are used in conjunction with categorical sampling schemes [paulus_rao_2021, gumbel_softmax, bengio_stochastic_2013]. In particular, a widely used stochastic scheme takes the following form, called Gumbel max:
@@
with $||$ denoting vector concatenation. Typically, however, neural networks interleave sequence processing layers with nonlinearly activated feed-forward networks [vaswani_attention_2017], as opposed to the affine transformation that we applied. We test the effect of a nonlinear readout, defined as
@@ Appendix F Related Work
LTI SSMs have been proposed as a more scalable alternative to Transformers that is also more performant on long sequences gu_hippo_2020, gu_combining_2021, gu_efficiently_2022, gu_parameterization_2022, gupta_diagonal_2022, smith_simplified_2023, orvieto_resurrecting_2023. A subsequent analysis showed promising language modeling performance with hybrid architectures based on gated SSMs and Attention ma2022mega, fu_hungry_2023. Time-variant hybrid systems were then shown to perform significantly better on practical tasks with the Mamba SSM which utilizes diagonal transition matrices gu_mamba_2023. Consequently, a line of work has demonstrated that hybrid models utilizing diagonal time-variant SSMs and attention exhibit better generalization to longer sequences, better general language modeling and reasoning performance [ren_2024_samba, de_griffin_2024, lenz_2025_jamba, wu2025transxssm, waleffe_2024_empirical].
@@
Concurrently, formal analyses of time-varying SSMs have shown limitations of the diagonal transition matrix structure merrill_illusion_2024, cirone_theoretical_2024, spurring a line of work investigating the utility of block-diagonal and unstructured matrices fan_advancing_2024, terzic2025sdssm, as well as more expressive structured transition matrices, most notably diagonal plus low-rank yang2024parallelizing, grazzi2024unlocking, peng_2025_rwkv, siems_2025_deltaproduct, with the original inspiration stemming from at least as far back as fast-weight programmers schlag2021linear, schlag2021learning, schmid_1992 whose computational structure contains products of generalized Householder matrices yang2024parallelizing, which enables the adoption of early algorithmic optimizations of such procedures wy_rep_1087. More recently, movahedi2025fixedpoint proposes fixed-point RNNs that can dynamically trade off expressivity with efficiency.
@@
As a lower bound on diagonal SSMs, sarrof2024expressivecapacitystatespace show that a stack of complex diagonal SSM layers can emulate any solvable automaton, with the stack depth proportional to the Krohn-Rhodes (KR) complexity of the transformation semigroup krohn_algebraic_1965, margolis_2024_decidability. As an upper bound on diagonal SSMs, merrill_illusion_2024 derives a bound on fixed-depth diagonal selective SSMs with logarithmic precision representation, placing them in the $\text{L-uniform }TC^{0}$ circuit complexity class. This complexity class encompasses solvable automata. DPLR matrices can emulate any automaton with depth proportional to KR-complexity and width linear in depth and arbitrary readout siems_2025_deltaproduct.
@@
Experimental analysis of sequential neural networks trained on formal language tasks goes back to at least [smith_zipser_1989, hochreiter_long_1997]. The exact experimental setup we build on is a more recent one, with the set of tasks taken from deletang_neural_2023, and the experimental setup borrowed from walker_2025_structured. All of the non-solvable group word problems were generated using modified sample generation code from liu_transformers_2023.